Find the maximum area of the ellipse x^2/a^2+y^2/b^2 which touches the line Y=3x+2

Grade:12

1 Answers

sai nath reddy
4 years ago
we know that c^=a^m^+b^ย  ย  ย  ย  ย  ย on comparing we get 4=9a^+b^ย  ย  ย  ย  ย  the values satisfies are a = 1/3 and b=square root threeย  ย then we know area of ellipse is piAB hence substitute them then we get area=pi/root threeย  ย ย